OpenStudy (anonymous):

find the angle between vector u: <-2, 5> and vector v:<-1, 3>

OpenStudy (jdoe0001):

well, first off get the cosine of the angle then get the arcCos of it \(\bf cos(\theta)=\cfrac{u \cdot v}{||u||\times ||v||}\)

OpenStudy (anonymous):

@jdoe0001 careful using \(\times\) there -- people may confuse it for a cross product

OpenStudy (jdoe0001):

hehe, yes it can get ambiguous

OpenStudy (jdoe0001):

I meant to leave it out, but then again some folks may wonder if it's a product or what

OpenStudy (jdoe0001):

\(\bf cos(\theta)=\cfrac{u \cdot v}{||u||\times ||v||} \implies \cfrac{\text{dot product}}{\text{product of magnitudes}}\)

OpenStudy (anonymous):

wait sorry... how do you find the magnitudes?

OpenStudy (anonymous):

@ses11 they are the norms of the vectors

OpenStudy (anonymous):

\[||u||=\sqrt{\sum_{k=1}^n (u_k)^2}\] for \(u=\langle u_1,...,u_n\rangle\).

OpenStudy (anonymous):

@oldrin.bataku so IIuII=\[\sqrt{-2^{2}+5^{2}}\]